OpenStudy (anonymous):

The area bounded by the curve y = x2 and the line y = 4 generates various solids of revolution when rotated as follows.

OpenStudy (anonymous):

it as to rotate it about the x and y axis

OpenStudy (anonymous):

for the x axis part i integrated from -2 to 2 of x^2

OpenStudy (anonymous):

for the y part should i integrate from 0 to 4 of sqrt y

OpenStudy (anonymous):

for y part INT pi * (sqrt y )^2) dy

OpenStudy (turingtest):

yes the part is\[V=\pi \int\limits_{0}^{4}(\sqrt{y})^2dy=\pi \int\limits_{0}^{4}ydy=(\pi/2)(16)=8 \pi\]

OpenStudy (turingtest):

about the y-axis I mean

OpenStudy (anonymous):

yep we agree on that

OpenStudy (turingtest):

How do we do the x part though? \[\pi \int\limits_{-2}^{2}4^2-x^2dx\]right?

OpenStudy (turingtest):

yes, that's it\[\pi \int\limits_{-2}^{2}4^2-x^2dx=\pi(16x-x^3/3)\]evaluated from -2 to 2 is\[\pi [(32-8/3)-(-32+8/3)]=\pi (64-16/3)=176 \pi/3\]

OpenStudy (anonymous):

wait are you using the washer method for the x part

OpenStudy (turingtest):

Yes. The outer radius is y=4 and the inner radius is x^2.

OpenStudy (anonymous):

got you

OpenStudy (anonymous):

what is it says to rotate about the line y=4